OpenStudy (anonymous):

light of wavelength 700nm is incident on a pair of slits, forming fringes 3.0mm apart on the screen. What is the fringe spacing when light of wavelength 350nm is used and the slit separation is doubled?

OpenStudy (anonymous):

Answers are a) 0.75mm b) 1.5mm c) 3.0mm d) 6.0mm

OpenStudy (anonymous):

I have a couple of equations in the book but I don't even know which one calculates fringe spacing... Eish! Appreciate any help :)

OpenStudy (sidsiddhartha):

ok are familiar with bragg's law which is ---- \[dsin \theta=n \lambda\]

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

Theta is the angle between the fringes, d the distance between the "lines" of the diffraction grating, n a whole number representing the order of the bright fringe, and lambda being the wavelength. Correct?

OpenStudy (sidsiddhartha):

yeah

OpenStudy (sidsiddhartha):

so \[\sin \theta=n \lambda/d\] so considering n=1 look lambda is halfed and slit separation d is doubled so \[\delta [\sin(\theta)]=\frac{ 1 }{ 4 }\] and it is generally spacing upon distence so fringe spacing when light of wavelength 350nm is used and the slit separation is doubled is=3/4=0.75 mm

OpenStudy (anonymous):

Okay two question more please: where does the three come from and so is sin theta equal to fringe spacing then?